# Feasibility of Learning Weighted Automata on a Semiring

### v3

@@ Abstract.
Since the seminal work by Angluin and the introduction of the $L^{*}$-algorithm, active learning of automata by membership and equivalence queries has been extensively studied to learn various extensions of automata. For weighted automata, algorithms for restricted cases have been developed in the literature, but so far there was no global approach or understanding how these algorithms could (or could not) apply in the general case. In this paper we chart the boundaries of the Angluin approach (using a class of hypothesis automata constructed from membership and equivalence queries) for learning weighted automata over a general semiring. We show precisely the theoretical limitations of this approach and classify functions with respect to how ‘guessable’ they are (corresponding to the existence and abundance of solutions of certain systems of equations). We provide a syntactic description of the boundary condition for a correct hypothesis of the prescribed form to exist. Of course, from an algorithmic standpoint, knowing that (many) solutions exist need not translate into an effective algorithm to find one; we conclude by relating our work to the existing literature with a discussion of some known conditions (and variants thereof) that suffice to ensure this, illustrating the ideas over several familiar semirings (including the natural numbers).
@@ 1. Introduction
Active learning has been extended to weighted automata but only for restricted classes: it has first been studied for fields [8, 7], and then principal ideal domains [21]. The Angluin framework has also been generalised to other extensions of automata and transducers [37, 38, 11, 9, 6, 5, 10, 16, 29, 23]. In [13] (and at first [22]), a categorical approach was proposed to encompass many of these cases. However, although this encompasses the case of weighted automata over fields, this framework does not include weighted automata in general. As often with weighted automata, finding a unified approach (working for all semirings) is not easy.
@@
In [21], van Heerdt, Kupke and Silva set a general framework for learning weighted automata, presenting a general learning algorithm, and focus on specifying two conditions on the semirings (ascending chain condition and progress measure), ensuring termination of the algorithm - this allows them to develop a learning algorithm à la Angluin for automata weighted on a principal ideal domain. Section 5 and Theorem 17 of [21] also present a specific example of an automaton (as well as a more general class) over the non-negative integers for which their algorithm does not terminate. In the present paper, we develop a general framework that gives a deeper understanding of the reasons for that.
Contribution:
@@
Our approach is somewhat different to what has been done so far: we look globally at all functions computed by finite state weighted automata and investigate whether they can potentially be learned by an Angluin-style algorithm. Working over an arbitrary semiring, we define levels of learnability and a hierarchy of functions with respect to these levels: the ones that definitely cannot be learned à la Angluin, the ones that might be, provided we are lucky or have a good algorithmic strategy, and the ones that definitely will be. Our main result in this regard is a syntactic condition on the functions that have the potential to be learned by an Angluin-style algorithm. For such a function there must exist a finite state automaton of a prescribed type computing it: we call these automata literal and they are closely related to residual automata. In cases already widely studied in the literature (e.g. where the weights are drawn from fields or integral domains), it turns out that every function computed by a finite state automaton can be computed by a literal automaton. Thus, in the existing literature, focus is (reasonably) on algorithmic issues. In this paper, we focus on picturing the landscape for active learning in arbitrary semirings, and showing the limitations; for instance over the non-negative integers, it is easily seen that the example of Heerdt, Kupke and Silva mentioned in the previous paragraph cannot be computed by a literal automaton. We do discuss a general algorithm (the one from [21]) and give some conditions on the target functions to ensure termination. In doing so, we refine the ascending chain condition and progress measure from [21] to pinpoint exactly what is needed, and we explain fully the reasons behind [21, Theorem 17].
@@
Within this framework, we then consider some properties of semirings that cause certain parts of our general hierarchy to collapse, and consider several specific well-studied examples of semirings (including the non-negative integers) in this context and provide a more detailed picture of the hierarchy for these. Note that the learning algorithm is based on equivalence queries, and one might ask whether the undecidability of equivalence for certain semirings [26] is problematic. Since in the application fields, equivalence can rarely be used as such (and often testing or approximation might be used instead), we believe having a framework that applies even in these cases is still of interest.
@@
The difficulty of this work was in setting a new framework and the (we believe) right notions for it. Once the correct definitions are in place, the results and proofs follow nicely. This new framework also opens avenues for research.
@@
In Section 2, we recall the definition of weighted automata, some specific examples, and outline the type of learning we are considering. We end this section with a set of questions that arise naturally, and which motivate the definitions and results in the rest of the paper. In Section 3, we explain what we mean by levels of learnability and define classes of functions depending on them. We give several characterisations for these functions, and relate our work back to the existing literature (specifically,  [21]), which focuses on algorithmic issues in specific settings. In Section 4, we picture the hierarchy based on the previous definitions, prove its strictness and give conditions which cause parts of the hierarchy to collapse. Throughout we raise a number of questions and indicate avenues for future progress, summarising these in 5.
@@ 2.1. Weighted automata and specific semirings
A finite alphabet is a finite set of symbols, called letters. A finite word is a finite sequence of letters and $\varepsilon$ denotes the empty word. We will also denote by $|w|$ the length of a word $w$ and by $|w|_{a}$ the number of occurrences of the letter $a$ in $w$.
A monoid $(M,\ast,e)$ is a set $M$ equipped with a binary operation $\ast$ that is associative and has a neutral element $e$. A semiring $(S,\oplus,\otimes,0_{S},1_{S})$ is a set $S$ equipped with two operations such that $(S,\oplus,0_{S})$ and $(S,\otimes,1_{S})$ are monoids, $\oplus$ is commutative, $\otimes$ distributes on the left and the right of $\oplus$ and $0_{S}$ is a zero for $\otimes$. A semiring is said to be commutative if $\otimes$ is commutative. Given a semiring $(S,\oplus,\otimes,0_{S},1_{S})$, two matrices $A$ and $B$ over $S$ can be multiplied as usual by $(AB)_{i,j}=\bigoplus_{k=1,\ldots,n}(A_{i,k}\otimes B_{k,j})$, provided the number, $n$, of columns of $A$ is equal to the number of rows of $B$. It is straightforward to check that this multiplication is associative.
@@
A weighted automaton $\mathcal{A}$ on a semiring $(S,\oplus,\otimes,0_{S},1_{S})$ over alphabet $\Sigma$ is defined by: a finite set of states $Q$; initial-state vector $\alpha\in S^{Q}$ (which we view as a row vector); for all letters $a$ of $\Sigma$, transition matrices $M(a)\in S^{Q\times Q}$; and final-state vector $\eta\in S^{Q}$ (which we view as a column vector). The automaton $\mathcal{A}$ computes a function $f_{\mathcal{A}}:\Sigma^{*}\rightarrow S$ via $f_{\mathcal{A}}(w_{1}\cdots w_{k})=\alpha M(w_{1})\cdots M(w_{k})\eta$. We say that $f_{\mathcal{A}}(w)$ is the value computed by $\mathcal{A}$ on input $w$.
A weighted automaton can be seen as a (labelled, directed) graph with set of vertices $Q$, such that (i) for each state $p\in Q$ there is an initial weight $\alpha_{p}$ and a final weight $\eta_{p}$; and (ii) for each pair of states $p,q\in Q$ and each $a\in\Sigma$ there is a transition (labelled edge) $p\xrightarrow{a\,:\,m}q$ where $m=M(a)_{p,q}$ is called the weight of the transition (and if $m=0_{S}$, then we often omit this edge from the graph). A run from a state $p$ to a state $q$ is a (directed) path in the graph, its weight is the product of $\alpha_{p}$ by the product of the weights on the transitions (taken in the order traversed) followed by $\eta_{q}$. The value computed on a word $w$ is the sum of all weights of runs labelled by $w$. A state $p$ such that $\alpha_{p}$ (resp. $\eta_{p}$) is not $0_{S}$ might be referred to as an initial (resp. final) state. We use both the matrix and graph vocabulary in this paper, as convenient.
Throughout the paper, $\Sigma$ denotes a finite alphabet and $S$ a semiring.
@@
Besides giving a general framework valid for any semiring, we consider several specific well-studied semirings that arise frequently in application areas. We will consider the semiring (,+,×$(\mathbb{R},+,\times,0,1)$ and its restriction to non-negative reals $\mathbb{R}_{\geq 0}$, integers $\mathbb{Z}$, and non-negative integers $\mathbb{N}$. We also consider the semiring ({}$(\mathbb{R}\cup\{-\infty\},\max,+,-\infty,0)$ and its restrictions to ${}\mathbb{Z}\cup\{-\infty\}$ and ${}\mathbb{N}\cup\{-\infty\}$, denoted by $\mathbb{R}_{\max}$, $\mathbb{Z}_{\max}$ and $\mathbb{N}_{\max}$ and the Boolean semiring $\mathbb{B}$. Note that all these semirings are commutative. Finally, we will consider the non-commutative semiring with domain the finite subsets of $\Sigma^{*}$, with operations given by union and concatenation ($XY=\{xy\mid x\in X,y\in Y\}$) and neutral elements $\emptyset$ and $\{\varepsilon\}$. This semiring will be denoted $\mathcal{P}_{\rm fin}(\Sigma^{*})$.
@@
We consider $\Sigma=\{a,b\}$. Figure 1 depicts several weighted automata where the initial (resp. final) states with weight $1_{S}$ are indicated with an ingoing (resp. outgoing) arrow, the other states have initial or final weights $0_{S}$. All six examples will be used in Section 4 and the example in Figure 1(a) will be used as a running example to illustrate the different notions we introduce. The automaton in Figure 1(a) is considered over $\mathbb{N}_{\max}$ and computes the number of $a$’s if the word starts with an $a$, the number of $b$’s if the word starts with a $b$, and $\infty$- ∞ on the empty word. The automaton depicted in Figure 1(b) will be considered over $\mathbb{R}_{\max}$ and computes the length of the longest block of consecutive $a$’s in a word. The automaton in Figure 1(c) is considered over $\mathbb{N}$ and on input $w$ computes: $2^{|w|_{a}}$ if $w$ starts with an $a$; $2^{|w|_{b}}$ if $w$ starts with a $b$; and 00 if $w$ is the empty word. The automaton from Figure 1(d) is viewed over $\mathbb{R}_{\geq 0}$ and computes $12^{n}-$ 1 on the words $a^{n}$ for all positive integers $n$, and 00 on any other word. Finally the automata in Figures 1(e) and 1(f) are over $\mathcal{P}_{\rm fin}(\Sigma^{*})$. The first one has outputs: $\emptyset$ on the empty word; $\{a^{|w|_{a}}\}$ on all words $w$ starting with an $a$; and $\{b^{|w|_{b}}\}$ on all words $w$ starting with a $b$. The second one outputs $\{a^{|w|_{a}},b^{|w|_{b}}\}$ on all input words $w$. It is worth mentioning here that several of the examples above could be viewed as examples over smaller semirings (e.g. Figure 1(b) could be considered over the semiring $\mathbb{N}_{\rm max}$. However, as we shall see below, changing the semiring over which one views a given example necessarily changes the space of hypothesis automata that one can construct using membership and equivalence queries.
@@
Since we are going to consider non-commutative semirings, care needs to be taken when talking about linear combinations. Given a set $Y$, an element $x$ of $S^{Y}$ and $\lambda$ in $S$, $\lambda\otimes x$ denotes the element of $S^{Y}$ computed from $x$ by multiplying every component of $x$ by $\lambda$ on the left. Given a subset $X\subseteq S^{Y}$, a left-linear combination over $X$ is one of the form $\bigoplus_{x\in X}\lambda_{x}\otimes x$ for some $\lambda_{x}$ in $S$, with only finitely many $\lambda_{x}$ different from $0_{S}$. The left-semimodule generated by $X$ is then the subset of $S^{Y}$ containing those elements that can be written as a left-linear combination over $X$. An element of the left-semimodule generated by $X$ will simply be said to be left-generated by $X$.
@@
Finally, we will use the mirror operation on functions $\Sigma^{*}\to S$. The mirror of a word $w=w_{1}w_{2}\cdots w_{n}$ where for all $i=1,\ldots,n$, $w_{i}\in\Sigma$ is the word $\bar{w}=w_{n}\cdots w_{2}w_{1}$. For $f:\Sigma^{*}\to S$, the mirror of $f$, denoted by $\bar{f}:\Sigma^{*}\to S$ is defined as $\bar{f}(w)=f(\bar{w})$. Given a weighted automaton $\mathcal{A}$, the mirror of $\mathcal{A}$ is obtained by reversing its transitions and swapping the initial-state vector and the final-state vector. Note that if $S$ is commutative and $f$ is computed by a weighted automaton $\mathcal{A}$ over $S$ then $\bar{f}$ is computed by the mirror of $\mathcal{A}$. However, there are cases when $S$ is non-commutative where a function can be computed by a weighted automaton over $S$ but its mirror cannot (for example, the function $w\mapsto\{w\}$ over $\mathcal{P}_{\rm fin}(\Sigma^{*})$ is computed by a weighted automaton over $\mathcal{P}_{\rm fin}(\Sigma^{*})$ but the mirror of this function ($w\mapsto\{\bar{w}\}$) is not).

### v4

@@ Abstract.
Since the seminal work by Angluin and the introduction of the $L^{*}$-algorithm, active learning of automata by membership and equivalence queries has been extensively studied to learn various extensions of automata. For weighted automata, algorithms for restricted cases have been developed in the literature, but so far there was no global approach or understanding how these algorithms could apply (or not) in the general case. In this paper we chart the boundaries of the Angluin approach. We use a class of hypothesis automata which are constructed, in Angluin’s style, by using membership and equivalence queries and solving certain finite systems of linear equations over the semiring, and we show the theoretical limitations of this approach. We classify functions with respect to how ‘guessable’ they are, corresponding to the existence of hypothesis automata computing a given function, and how such an hypothesis automaton can be found. Of course, from an algorithmic standpoint, knowing that a solution (hypothesis automaton) exists need not translate into an effective algorithm to find one. We relate our work to the existing literature with a discussion of some known properties ensuring algorithmic solutions, illustrating the ideas over several familiar semirings (including the natural numbers).
@@ 1. Introduction
Active learning has been extended to weighted automata but only for restricted classes: it has first been studied for fields [9, 7], and then principal ideal domains [36]. The Angluin framework has also been generalised to other extensions of automata and transducers [40, 41, 12, 10, 6, 5, 11, 18, 28, 38]. In [15] (see also [37] [8] and [34] for earlier works), a categorical approach was proposed to encompass many of these cases. However, although this encompasses the case of weighted automata over fields, this framework does not include weighted automata in general. As often with weighted automata, finding a unified approach (working for all semirings) is not easy.
@@
In [36], van Heerdt, Kupke, Rot and Silva set a general framework for learning weighted automata, present a general learning algorithm, and focus on specifying two conditions on the semirings (ascending chain condition and progress measure), ensuring termination of the algorithm - this allows them to develop a learning algorithm à la Angluin for automata weighted on a principal ideal domain. More recently, a polynomial-time algorithm within this general framework has been developed in [13] for the principal ideal domain $\mathbb{Z}$. Section 5 and Theorem 17 of [36] also present a specific example of an automaton (as well as a more general class) over the non-negative integers for which their algorithm does not terminate. In the present paper, we develop a general framework that gives a deeper understanding of the reasons for that.
Contributions:
@@
Our approach is somewhat different to what has been done so far: we look globally at all functions computed by finite state weighted automata and investigate, for each of them, whether they can potentially be learned by an Angluin-style algorithm, i.e. where the automata that are guessed (hypothesis automata, see Definition 2.10) are constructed in a prescribed way using finitely many membership queries (to be explained in detail below). Working over an arbitrary semiring, we define levels of guessability and a hierarchy of functions with respect to these levels: from the ones for which no such guess will be correct to the ones where there is a simple theoretical strategy to construct a correct guess. Our main result in this regard (Theorem 3.6) is a syntactic condition on the functions that have the potential to be learnt. For such a function (weakly guessable functions, see Definition 3.1) there must exist a finite state automaton of a prescribed type computing it: we call these automata literal (Defintion 3.2) and they are closely related to residual automata. In cases already widely studied in the literature (e.g. where the weights are drawn from fields or integral domains), it turns out that every function computed by a finite state automaton can be computed by a literal automaton. Thus, in the existing literature, focus is (reasonably) on algorithmic issues. In this paper, we focus on picturing the landscape for active learning in arbitrary semirings, and showing the limitations; for instance over the non-negative integers, it is easily seen that the example of van Heerdt, Kupke, Rot and Silva mentioned in the previous paragraph cannot be computed by a literal automaton.
@@
Within this framework, we then consider some properties of semirings that cause certain parts of our general hierarchy to collapse, and consider several specific well-studied examples of semirings (including the non-negative integers) in this context and provide a more detailed picture of the hierarchy for these (Section 4). Note that the learning process is based on equivalence queries, and one might ask whether the undecidability of equivalence for certain semirings [25] is problematic. Since in the application fields, equivalence can rarely be used as such (and often testing or approximation might be used instead), we believe having a framework that applies even in these cases is still of interest.
@@
The difficulty of this work was in setting a new framework and the (we believe) right notions for it. Once the correct definitions are in place, the results and proofs follow nicely.
@@
In Section 2, we recall the definition of weighted automata, some specific examples, and outline the type of learning we are considering. We end this section with a set of questions that arise naturally, and which motivate the definitions and results in the rest of the paper. In Section 3, we explain what we mean by levels of guessability and define classes of functions depending on them. We give several characterisations for these classes of functions, and relate our work back to the existing literature (specifically,  [36]), which focuses on algorithmic issues in specific settings. In Section 4, we picture the hierarchy based on the previous definitions, prove its strictness and give conditions which cause parts of the hierarchy to collapse. Throughout we raise a number of questions and indicate directions for future progress, summarising these in Section 5.
@@ 2.1. Weighted automata and specific semirings
We begin by recording some preliminary definitions and notation. A finite alphabet $\Sigma$ is a finite set of symbols, called letters. A finite word is a finite sequence of letters and $\varepsilon$ denotes the empty word. We write $\Sigma^{*}$ to denote the set of all words over the alphabet $\Sigma$, including the empty word. We will also denote by $|w|$ the length of a word $w$ and by $|w|_{a}$ the number of occurrences of the letter $a$ in $w$.
A monoid $(M,\ast,e)$ is a set $M$ equipped with a binary operation $\ast$ that is associative and has a neutral element $e$. For example, the set $\Sigma^{*}$ together with the binary operation of concatenation of words and the empty word $\varepsilon$ is a monoid; it is the free monoid on the set $\Sigma$. A semiring $(S,\oplus,\otimes,0_{S},1_{S})$ is a set $S$ equipped with two operations such that $(S,\oplus,0_{S})$ and $(S,\otimes,1_{S})$ are monoids, $\oplus$ is commutative, $\otimes$ distributes on the left and the right of $\oplus$ and $0_{S}$ is a zero for $\otimes$. A semiring is said to be commutative if $\otimes$ is commutative.
Recall that for a set $Y$, we write $S^{Y}$ to denote the set of all functions from $Y$ to $S$. For non-empty (but possibly infinite) countable sets $I$ and $J$ we say that $A\in S^{I\times J}$ is a matrix, with rows indexed by the set $I$ and columns indexed by the set $J$. For $A\in S^{I\times J}$ and $B\in S^{J\times K}$ where $J$ is finite we define the product $AB$ in the usual way via $(AB)_{i,k}=\bigoplus_{j\in J}(A_{i,j}\otimes B_{j,k})$. Analogously, it will sometimes also be convenient to view elements of $S^{J}$ as (row or column) ‘vectors’ with entries indexed by $J$; for $\underline{x}\in S^{J}$ and $j\in J$ we write $\underline{x}_{j}$ to denote the value computed by the function $\underline{x}$ on input $j$, and if $J$ is finite we define the product $A\underline{x}$ where $A\in S^{I\times J}$ in the obvious way. It is also straightforward to check that this multiplication is associative, that is, for all $A\in S^{I\times J}$, $B\in S^{J\times K}$, $C\in S^{K\times L}$, where $J$ and $L$ are both finite, we have $A(BC)=(AB)C$.
@@
A weighted automaton $\mathcal{A}$ on a semiring $(S,\oplus,\otimes,0_{S},1_{S})$ over alphabet $\Sigma$ is defined by: a finite set of states $Q$; initial-state vector $\underline{\alpha}\in S^{Q}$ (which we view as a row vector); for all letters $a$ of $\Sigma$, transition matrices $M(a)\in S^{Q\times Q}$; and final-state vector $\underline{\eta}\in S^{Q}$ (which we view as a column vector). The automaton $\mathcal{A}$ computes a function $f_{\mathcal{A}}:\Sigma^{*}\rightarrow S$ via $f_{\mathcal{A}}(w_{1}\cdots w_{k})=\underline{\alpha}M(w_{1})\cdots M(w_{k})\underline{\eta}$. We say that $f_{\mathcal{A}}(w)$ is the value computed by $\mathcal{A}$ on input $w$.
A weighted automaton can be seen as a (labelled, directed) graph with set of vertices $Q$, such that (i) for each state $p\in Q$ there is an initial weight $\underline{\alpha}_{p}$ (represented by an incoming arrow at state $p$ with label $\underline{\alpha}_{p}$) and a final weight $\underline{\eta}_{p}$ (represented by an outgoing arrow at state $p$ with label $\underline{\eta}_{p}$); and (ii) for each pair of states $p,q\in Q$ and each $a\in\Sigma$ there is a transition (labelled edge) $p\xrightarrow{a\,:\,m}q$ where $m=M(a)_{p,q}$ is called the weight of the transition (and if $m=0_{S}$, then we often omit this edge from the graph). A run from a state $p$ to a state $q$ is a (directed) path in the graph and the weight of a run from $p$ to $q$ is the product of $\underline{\alpha}_{p}$ by the product of the weights on the transitions (taken in the order traversed) followed by $\underline{\eta}_{q}$. The value computed on a word $w$ is the sum of all weights of runs labelled by $w$. A state $p$ such that $\underline{\alpha}_{p}$ (resp. $\underline{\eta}_{p}$) is not $0_{S}$ is referred to as an initial (resp. final) state. We use both the matrix and graph vocabulary in this paper, as convenient.
Notations.
Throughout the paper, $\Sigma$ denotes a finite alphabet and $S$ a semiring (which will not be assumed to be commutative unless explicitly stated). The lower case symbols $a,b,c$ will be exclusively used to denote letters of an alphabet, whilst the symbols $u,v,w$ will be exclusively used to denote words over the alphabet $\Sigma$. We use calligraphic upper-case symbols such as $\mathcal{A},\mathcal{B},\mathcal{H}$ to denote weighted automata, whilst the lower-case symbols $f,g,h$ are always used to denote functions from $\Sigma^{*}$ to $S$. Greek letters such as $\alpha,\beta,\lambda,\mu,\rho,\eta$ will be used to denote elements of the semiring $S$, upper-case symbols $A,B,\ldots,F$ will be used to denote (possibly infinite) matrices over $S$, and underlined symbols (e.g. $\underline{x}$ or $\underline{\lambda}$) will denote (possibly infinite) vectors over $S$; in both cases the index sets used vary but will be made explicit. Upper-case symbols $I,J,\ldots,Z$ will typically be used for sets; frequently these will be (countable) subsets of $S^{\Sigma^{*}}$ or $\Sigma^{*}$, and moreover such subsets will frequently be used as indexing sets.
@@
Besides giving a general framework valid for any semiring, we consider several specific well-studied semirings that arise frequently in application areas. We will consider the semiring (,+,×$(\mathbb{R},+,\times,0,1)$ and its restriction to non-negative reals $\mathbb{R}_{\geq 0}$, integers $\mathbb{Z}$, and non-negative integers $\mathbb{N}$. We also consider the semiring ({}$(\mathbb{R}\cup\{-\infty\},\max,+,-\infty,0)$ and its restrictions to ${}\mathbb{Z}\cup\{-\infty\}$ and ${}\mathbb{N}\cup\{-\infty\}$, denoted by $\mathbb{R}_{\max}$, $\mathbb{Z}_{\max}$ and $\mathbb{N}_{\max}$ and the Boolean semiring $\mathbb{B}$. Note that all these semirings are commutative. Finally, we will consider the non-commutative semiring with domain the finite subsets of $\Sigma^{*}$, with operations given by union (playing the role of addition) and concatenation ($XY=\{xy\mid x\in X,y\in Y\}$, playing the role of multiplication) and neutral elements $\emptyset$ and $\{\varepsilon\}$. This semiring will be denoted $\mathcal{P}_{\rm fin}(\Sigma^{*})$.
@@
We consider $\Sigma=\{a,b\}$. Figure 1 depicts several weighted automata where the initial (resp. final) states with weight $1_{S}$ are indicated with an ingoing (resp. outgoing) arrow, the other states have initial or final weights $0_{S}$. All six examples will be used in Section 4 and the example in Figure 1(a) will be used as a running example to illustrate the different notions we introduce. The automaton in Figure 1(a) is considered over $\mathbb{N}_{\max}$ and computes the number of $a$’s if the word starts with an $a$, the number of $b$’s if the word starts with a $b$, and $\infty$- ∞ on the empty word. The automaton depicted in Figure 1(b) will be considered over $\mathbb{R}_{\max}$ and computes the length of the longest block of consecutive $a$’s in a word. The automaton in Figure 1(c) is considered over $\mathbb{N}$ and on input $w$ computes: $2^{|w|_{a}}$ if $w$ starts with an $a$; $2^{|w|_{b}}$ if $w$ starts with a $b$; and 00 if $w$ is the empty word. The automaton from Figure 1(d) is viewed over $\mathbb{R}_{\geq 0}$ and computes $12^{n}-$ 1 on the words $a^{n}$ for all positive integers $n$, and 00 on any other word. Finally the automata in Figures 1(e) and 1(f) are over $\mathcal{P}_{\rm fin}(\Sigma^{*})$. The first one has outputs: $\emptyset$ on the empty word; $\{a^{|w|_{a}}\}$ on all words $w$ starting with an $a$; and $\{b^{|w|_{b}}\}$ on all words $w$ starting with a $b$. The second one outputs $\{a^{|w|_{a}},b^{|w|_{b}}\}$ on all input words $w$. It is worth mentioning here that several of the examples above could be viewed as examples over smaller semirings (e.g. Figure 1(b) could be considered over the semiring $\mathbb{N}_{\rm max}$). However, as we shall see below, changing the semiring over which one views a given example necessarily changes the space of hypothesis automata that one can construct using membership and equivalence queries.
@@
Since we are going to consider non-commutative semirings, care needs to be taken when talking about linear combinations. Given a set $Y$, $S^{Y}$ is a left $S$-module via the following action: for each element $\underline{x}$ of $S^{Y}$ and $\lambda$ in $S$, $\lambda\otimes\underline{x}$ denotes the element of $S^{Y}$ computed from $\underline{x}$ by multiplying every component of $\underline{x}$ by $\lambda$ on the left. Given a subset $X\subseteq S^{Y}$, a left-linear combination over $X$ is one of the form $\bigoplus_{\underline{x}\in X}\lambda_{\underline{x}}\otimes\underline{x}$ for some $\lambda_{\underline{x}}$ in $S$, with only finitely many $\lambda_{\underline{x}}$ different from $0_{S}$. The left-semimodule generated by $X$ is then the subset of $S^{Y}$ containing those elements that can be written as a left-linear combination over $X$. An element of the left-semimodule generated by $X$ will simply be said to be left-generated by $X$. (Dually, one can view $S^{Y}$ as a right $S$-module, and make the corresponding definitions of right-linear combination, right-semimodule generated by $X$ etc.)
@@
Finally, we will use the mirror operation on functions $\Sigma^{*}\to S$. The mirror of a word $w=w_{1}w_{2}\cdots w_{n}$ where for all $i=1,\ldots,n$, $w_{i}\in\Sigma$ is the word $\bar{w}=w_{n}\cdots w_{2}w_{1}$. For $f:\Sigma^{*}\to S$, the mirror of $f$, denoted by $\bar{f}:\Sigma^{*}\to S$ is defined as $\bar{f}(w)=f(\bar{w})$. Given a weighted automaton $\mathcal{A}$, the mirror of $\mathcal{A}$ is obtained by reversing each of its transitions and swapping the initial-state vector and the final-state vector. For example, the automaton in Figure 1B is equal to its mirror. Note that if $S$ is commutative and $f$ is computed by a weighted automaton $\mathcal{A}$ over $S$ then $\bar{f}$ is computed by the mirror of $\mathcal{A}$. However, there are cases when $S$ is non-commutative where a function can be computed by a weighted automaton over $S$ but its mirror cannot (for example, the function $w\mapsto\{w\}$ over $\mathcal{P}_{\rm fin}(\Sigma^{*})$ is computed by a weighted automaton over $\mathcal{P}_{\rm fin}(\Sigma^{*})$ but the mirror of this function ($w\mapsto\{\bar{w}\}$) is not).